OpenStudy (anonymous):

Write an equation of a circle with the given center and radius. center (–7, –6) and radius 2 Help me please?

OpenStudy (anonymous):

Use the equation \[(x-h)^2+(y-k)^2=r^2\] where h and k are the x and y coordinates respectively, and r is the radius.

OpenStudy (anonymous):

so (x+7)^2+(y+6)^2=2^2

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Can you help me with another one!

OpenStudy (anonymous):

? please

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Graph 4x2 + y2 = 9. What are its lines of symmetry?

OpenStudy (anonymous):

4x^2+y^2=9.

OpenStudy (anonymous):

@yaro99

OpenStudy (anonymous):

the equation of an ellipse is \[(x-h)^2/a^2+(y-k)^2/b^2=1\] divide your equation by 9 so that you have it in the above form.  (This is a random ellipse, just to show what the a and b values are, yours may or may not look like it) Plot the equation so you see what it looks like, then find the vertical and horizontal lines that divide the circle in two:  The lines of symmetry are in the form x=h and y=k, in this case h and k are both 0, meaning that the ellipse is centered at the origin and you have x=0 and y=0 (you could skip straight to this step, but graphing makes it easier to visualise)